A cyclist is traveling 85 miles in 5 hours against the wind. He travels 174 miles in 6 hours with the wind. What is the rate of the still air and the rate of the wind.
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

D=RT
85=(R-W)*5 AGAINST THE WIND.
85=5R-5W
5R=85+5W
R=(85+5W)/5
R=17+W
174=(R+W)6 WITH THE WIND.
174=6R+6W
6R=174-6W
R=(174-6W)/6
R=29-W
17+W=29-W
2W=29-17
2W=12
W=12/2
W=6 MPH.
R=17+6=23 THE SPEED OF THE BIKE IN STILL AIR.
R=29-6=23
PROOF:
85=(23-6)5
85=17*5
85=85
174=(23+6)*6
174=29*6
174=174


Answer by Alan3354(69443)

What does the wind have to do with the speed of a bicycle?
Have you ever ridden a bicycle in the wind?
-------------
The best answer is, "We don't know."